# Online Navigation Planning for Long-term Autonomous Operation of Underwater Gliders

### v1

@@ V-E System Components and Integration
Glider Long-term Autonomy Engine. The GLAE system contributed in this paper implements the methodology described in Section IV and is set up as a service deployed to on-premise computational infrastructure. It is capable of controlling several gliders simultaneously. GLAE fetches the most recent current forecasts from Copernicus or the Met Office Marine Data Service daily once they are published to inform navigation planning.
C2 Communications Backbone. This open-source component [38] provides publish-subscribe functionality. It is powered by a RabbitMQ message queue. A variety of clients can read and write messages in a suite of standardised formats [39] via HTTP APIs that are access-controlled. In the remainder of this section, we indicate specific message formats from this repository using monospace font (e.g., platform_status).
SFMC Adapter. The Slocum Fleet Mission Control (SFMC) Adapter is a microservice that converts between the standardised formats of the Communications Backbone and the specific file formats required by Slocum gliders.
@@
SFMC. This proprietary component is the software suite created by the Slocum glider manufacturer. It provides interfaces for controlling the glider, either manually via a terminal or programmatically using APIs. It enables transferring files containing waypoints and control parameters to the glider by placing them in a dedicated directory from which they are read. SFMC communicates with gliders directly via an Iridium satellite link.
@@
C2 Piloting User Interface. This web-based interface can be used to configure missions for, monitor, and pilot marine autonomous systems. It is complementary to the SFMC web-based interfaces and unifies piloting interfaces for a variety of platforms beyond Slocum gliders [37].
V-F Workflows
Next, we describe the two main workflows. The Mission Configuration Workflow is higher-level and executed once at the start of a mission. The Online Planning Workflow is lower-level and executed after every glider surfacing in order to provide up-to-date control instructions so as to best achieve the mission goals.
@@ V-F Workflows
Mission Configuration Workflow. The workflow to configure a glider mission involves the use of the C2 Piloting UI to indicate the glider for the mission, the list of goals $\mathbf{p}_{\text{goal}}^{(i)}]$, and permitted instruction sets $\mathcal{K}$. A planning_configuration message is generated and published to the Backbone. The GLAE system, upon receiving the message, initialises the tracking information in persistent storage. From this point onwards, GLAE will track the mission goals and respond to post-surfacing updates from the glider with navigation plans.
Online Planning Workflow. This workflow achieves the full glider control loop. The steps are illustrated in Figure 6, wherein the numbers above the arrows match the bullet point numbering below.
  1. 1.
    Upon surfacing, the glider calls in to SFMC. Over satellite, it relays status information (time, GPS fix) and transfers scientific data while awaiting instructions as configured by its mission script. The instructions will be received at the last step of this workflow.
  2. 2.
    The SFMC Adapter converts the status information and publishes a platform_status message to the Communications Backbone message queue.
  3. 3.
    The platform_status message is delivered to GLAE. The system creates a corresponding problem instance $I$.
  4. 4.
    The Planner (Algorithm 2) is invoked using the problem instance $I$.
  5. 5.
    The Simulator (Algorithm 1) is repeatedly called by the Planner. A number of dives on the order of $10^{3}-10^{4}$ is simulated by each thread of the Planner. This is the most computationally intensive step, requiring several minutes to complete.
  6. 6.
    The best action $\mathbf{a}$ is determined by the Planner and converted to a list of waypoints using Algorithm 3. Corresponding mission_plan (the message format representing a waypoint list) and instruction_set messages are generated and published to the Communications Backbone.
  7. 7.
    The messages are processed by the SFMC Adapter and placed as files in the appropriate directory on SFMC.
  8. 8.
    The files are transmitted to the glider and the corresponding parameters are used to execute the next dive.
@@
It is worth reiterating that steps 1 and 8, which involve communication over satellite, are not fully reliable. In preliminary deployments, the rate of successful transfer of waypoint lists and instruction sets was between 88% and 90% overall, with further significant drops during periods with adverse weather conditions. Given that navigation plans cover a longer horizon than individual dives (on the order of a day rather than hours), the glider was able to continue its mission.
VI Experimental Setup
VI-A Deployments
@@
Our system was evaluated in two campaigns in the North Sea: MOGli (one glider) and eSWEETS3 (two gliders simultaneously). In these deployments, gliders were operated by NOC on behalf of external partners Met Office and the University of East Anglia. A transect formed of two goal positions was used in all deployments. Autonomous control by GLAE was used through the workflows described in the previous section to repeatedly navigate the transects. Deployment coordinates and summary statistics are given in Tables I and II. A map of the three field deployments is shown in Figure 7.
TABLE I: Deployment goal locations and parameters. All distances are given in km.
@@
Figure 7: Locations of the field deployments in the North Sea.
MOGli. Met Office Glider (MOGli) is an ongoing joint project between the UK Met Office (MO) and NOC. Measurements of ocean properties are gathered by a single glider operated by NOC and transmitted periodically to the MO. This data is used to refine MO models. The use of GLAE to pilot the glider leads to a virtuous circle as AMM15 forecasts are used to derive navigation plans, while the scientific data itself is used to improve the forecasts.
For this deployment, a region that exhibits low bottom temperature variance was selected in accordance with the MOGli scientific objectives. Fully autonomous control operated from 2 May to 3 July 2025, following preliminary experiments that began in November 2024 and built towards achieving the full glider control loop. A pause in autonomous control was in place from 21 May to 7 June 2025 while the glider was recovered, recharged, and redeployed to the experiment region.
The valid action space allowed relative bearings $\alpha\in\{-40,-20,0,20,40\}$. A value of $n_{\text{yos}}=5$ yos was chosen to lead to an operationally appropriate dive duration (approximately $3$ hours). A target depth of $z_{\text{bottom}}=150$ was used, which exceeds the ocean floor. This was driven by the requirement to sample the entire water column, and hence the glider uses the altimeter to inflect when close to the ocean floor, a feature also supported by our simulator.
eSWEETS3. In the enabling Sustainable Wind Energy Expansion in Seasonally Stratified Seas (eSWEETS3) project, two Slocum gliders were deployed around the Kincardine, UK offshore wind farm. The gliders were positioned upstream (referred to as eSWEETS3 North) and downstream (eSWEETS3 South) of the wind farm to assess the impacts of the turbine structures on the vertical mixing in the area, which influences oxygen availability, nutrient supplies, and biological productivity.
The project deployments were performed in two phases. In the first phase (May - August 2025), gliders were piloted manually using standard procedures. In the second phase (August - October 2025), GLAE was used to control the gliders following preliminary experiments and extensions to system functionality. Fully autonomous operation was in place from 2 September to 30 October 2025 for eSWEETS3 North and from 18 September to 30 October 2025 for eSWEETS3 South.
The valid action space allowed relative bearings $\alpha\in\{-90,-60,-30,0,30,60,90\}$, a wider range than in MOGli to account for the strong effect of tidal currents. $n_{\text{yos}}=10$ yos were used, leading to dive durations of approximately $4$ hours in this shallower region. A target depth of $z_{\text{bottom}}=150$ was used which, as in MOGli, exceeds the bathymetric limit.
TABLE II: Deployment summary statistics.
@@
The eSWEETS3 deployments were subject to additional operational requirements. As the gliders were deployed in areas with strong tidal currents, there was a risk of vehicle loss if drifting away from the mission area, especially inland. Safety polygons were defined and additional functionality was implemented such that, if the glider drifts outside of the safety polygon, Algorithm 3 is replaced with a simplified navigation plan aiming directly for the center of the polygon. Normal navigation was resumed upon re-entry. Furthermore, as continuing to sample along the transect upon re-entry was deemed more important for the scientific objective than turning back to achieve the goal, the system switched to the next goal if the glider had already progressed past 90% of the transect before exiting the safety polygon.
Transect integrity. Occasionally, navigation plans were not generated or followed by the glider due to planned (e.g., software upgrades) or unplanned (e.g., anomalies causing abort of the mission) concerns. Even though these events typically impacted a single surfacing, they compromise the transect validity. Such transects were excluded from the final results.
VI-B Simulation Experiments
We present a set of experiments in simulation whose setup mirrors the deployments listed above. They are an idealised version of the real-world experiments, wherein file transfers always succeed and currents experienced are well-represented by the forecasts. Underwater goal achievement detection (Section V-C) and the additional functionality for eSWEETS3 (Section VI-A) are not supported in simulation. The simulator and planner parameters are otherwise set identically to those used in the field deployments.
We select three scenarios for each deployment to be replayed in simulation. A scenario consists of the starting state (glider position and time) and the goal assigned to the glider in the real world. Scenarios are used as starting conditions for the simulation. To determine the scenarios, we first assess the number of dives required by the Straight-to-Goal method (described in the next subsection) to complete each transect averaged over $200$ repetitions. We pick the starting conditions for the transects with the minimum (Favourable), median (Neutral), and maximum (Unfavourable) dives required, on average, for the baseline to reach the goal. This yields scenarios that represent a range of conditions experienced by the glider during the deployment periods. We then re-run both the Planner and Straight-to-Goal over $1000$ random seeds to report the final results.
VI-C Baseline
As a baseline for the Planner, we consider the Straight-to-Goal approach, which always selects the action with $\alpha=0$. It is used for comparisons in both the simulated and real-world deployments. For an equitable comparison to the Planner, when Straight-to-Goal is used in the real-world deployments, the glider remains under the autonomous control of GLAE.
@@ VI-C Baseline
In the real-world deployments, only one of the two methods can control the glider at any one time. An alternating-control approach was taken, in which Planner and Straight-to-Goal took turns to control the glider after the achievement of $2$ goals (equivalently, the algorithm was switched after every back-and-forth transect traversal). Even though the currents experienced by the two algorithms in the real world differ, we expect that any bias thus introduced is removed in expectation. Furthermore, in simulated experiments, we may use the exact same currents when evaluating both methods and replay the same scenarios as many times as desired with different realisations of the uncertainty.
@@
For MOGli, alternating-control was followed throughout. For eSWEETS3, exclusive use of Straight-to-Goal was mandated initially by operational concerns. GLAE was subsequently set to issue Planner waypoints for a comparable duration. For the final period, the algorithms were switched after every transect traversal as in MOGli.
VI-D Metrics
@@
Planner. To evaluate the planner, we report the total transect duration (i.e., the cumulative cost as defined in Section IV-B), which is the objective that the method optimises for. As secondary metrics, we report the number of dives (highly correlated with total duration) and the total distance covered for each traversal. All ± entries in tables denote 95% confidence intervals.
@@
Simulator. To evaluate the simulator, we report the true and simulated dive distances and durations, as well as their differences (deltas). These metrics are assessed on the dives recorded during the live deployment, which were not used for simulator parameter fitting or validation. Furthermore, as the simulator is stochastic and produces several plausible surfacing locations and times, we report statistics with respect to the sample means obtained by drawing $40$ samples for each dive.
VI-E Simulator and Planner Parameter Optimisation
@@
To optimise the simulator parameters (Section IV-D), we used historical glider states / actions and current forecasts gathered in preliminary deployments (concretely, February - March 2025 for MOGli and May - June for eSWEETS3). We use the files generated by the glider and archived AMM15 forecasts for this purpose. However, such data can also be obtained publicly through the British Oceanographic Data Centre (BODC) [40] and Copernicus [41], with which GLAE is compatible.
@@
The dives are split randomly into training and validation sets. Parameters are optimised using Bayesian optimisation (BO) on the training set, and the best-performing parameters on the validation set with respect to $J(\Theta)$ are selected. We apply several data cleaning steps by excluding dives that did not terminate due to normal surfacing, those with speed $2$ standard deviations outside the mean, and with a difference between heading and surface location greater than a $75$ degree threshold. $n_{\text{SIM-BO}}=500$ iterations of BO were used as the scoring function plateaued before this step in all experiments.
@@
In terms of planner parameters, we use $n_{\text{threads}}=8$ parallel search trees for both deployments. We set $n_{\text{trials}}=10000$ and $n_{\text{trials}}=5000$ for MOGli and eSWEETS3 respectively, as for the latter deployment less data is transmitted and therefore less online planning time is available. These parameters are dictated externally by the available resources. We found that the planner hyperparameters (exploration constant, progressive widening parameters, maximum tree depth) are sensitive to simulator parameters. Therefore, after fixing simulator parameters, planner hyperparameters are optimised downstream, also with BO.
VII Results
VII-A Simulator Evaluation
@@
Figure 8 shows the progress of the scoring function $J(\Theta)$ during the steps of the Bayesian optimisation procedure on the MOGli data. Several values for the spread penalty $\lambda_{\text{reg}}$ were tested, and the parameters that obtained the best $J(\Theta)$ on the validation set were chosen ($\lambda_{\text{reg}}=1$ for MOGli, corresponding to the figure shown and $\lambda_{\text{reg}}=0$ for eSWEETS3).
In Table III, we report several metrics concerning simulator performance on the deployment data (not seen during training or model selection). The distances of the dives output by the simulator were close to their true average on eSWEETS3, but noticeably underestimated for MOGli. The distance between the true and simulated surfacing locations is around $1.5$ km on average for all deployments (though it is inherently challenging to draw conclusions by comparing an individual sample with the empirical mean of the simulator distribution). In contrast, the dive durations were more accurate for MOGli, as there is increased variation in the bathymetry between the eastern and western ends of the eSWEETS3 region.
We note that there are many possible causes for predictive errors in the simulator, which include: forecasts being wrong, the steady-state approximation of our simulator, mismatches between the physical glider used for simulator optimisation and the one used in the real world (distinct vehicles with variations in payload and calibration), and noise in how the glider executes the control commands. It is highly likely that all these factors are at play and, while simulator accuracy can certainly be improved, some degree of error is unavoidable.
TABLE III: Comparison of simulated and true dive metrics averaged over all dives executed in the real-world deployments.
@@
Simulated dive distance (m) 2119 5586 4974
Distance delta (m) 1448 1735 1396
@@
Duration delta (s) 345 1721 1539
VII-B Planner Evaluation in Simulation
TABLE IV: Results comparing the Straight-to-Goal and Planner algorithms across several scenarios in simulation. Bold indicates the better-performing algorithm.
Deployment Scenario Algorithm Mean Duration (hrs) Mean Number Dives Mean Transect Length (km)

### v2

@@ 5.C System Components and Integration
The GALE system implements the methodology described in Section 4 and is a service deployed to an on-premise server. It is capable of controlling several gliders simultaneously and fetches the most recent current forecasts daily. In this section, we describe the components that are required besides GALE to achieve the full control loop of the glider.
C2 Communications Backbone. This open-source component [40, 41] developed by the UK National Oceanography Centre (NOC) enables real-time transfer of vehicle information and control instructions using standardised APIs [42].
@@
SFMC. Slocum Fleet Mission Control (SFMC) provides interfaces for controlling gliders manually or programmatically. It enables transferring files containing waypoints and control parameters by placing them in a dedicated directory from which they are transmitted via an Iridium satellite link.
@@
C2 Piloting User Interface. This web-based interface supports configuring missions for, monitoring, and piloting marine autonomous systems. It is complementary to SFMC and unifies piloting interfaces for several platforms [40].
5.D Workflows
Two workflows are illustrated in Figure 8.
@@ 5.D Workflows
Mission Configuration Workflow. This workflow is executed once at the start of a mission. It involves the use of the C2 Piloting UI to indicate the glider, the list of goals $\{\mathbf{p}_{\text{goal}}^{(i)}\}$, and permitted instruction sets $\mathcal{K}$. The GALE system initialises the tracking information in persistent storage. From this point onwards, GALE will track the mission goals and respond to post-surfacing updates from the glider with navigation plans.
Online Planning Workflow. This workflow achieves the full glider control loop. Upon surfacing, the glider calls in to SFMC (1). Over satellite, it relays status information (time, GPS fix) and transfers scientific data while awaiting instructions from GALE, which creates a corresponding problem instance $I$ (2, 3). The Planner (Algorithm 2) is invoked using $I$, which calls the Simulator (Algorithm 1) on the order of $10^{3}-10^{4}$ times in each thread (4,5). The best action $\mathbf{a}$ is determined by the Planner and converted to a waypoint list and instruction set, which are published to the Backbone (6). The messages are placed as files in the appropriate directory on SFMC (7). They are transmitted to the glider and are used to execute the next dive (8).
@@
It is worth reiterating that steps 1 and 8, which involve communication over satellite, are not fully reliable. In preliminary deployments, the rate of successful transfer of waypoint lists and instruction sets was approximately 89%, with further significant drops during periods with adverse weather conditions. Since navigation plans cover a longer horizon than individual dives (on the order of a day rather than hours), the glider was able to continue its mission.
5.E System Configuration
Configuring the system for a new deployment requires the following steps:
  • The scientific stakeholders supply the mission goals $\{\mathbf{p}_{\text{goal}}^{(i)}\}$ and the goal tolerance radius $\rho$.
  • The glider pilots specify the permitted instruction sets $\mathcal{K}$ that are appropriate for the chosen vehicle and the scientific objectives.
  • Relevant data from historical glider deployments is identified and retrieved (e.g., from public repositories [43]). Alignment to the upcoming deployment in terms of location, vehicle type, and time period are important for the faithfulness of the simulator.
  • The GALE system administrator performs simulator parameter tuning (Section 4.D) using the historical data. Planner parameter tuning is also performed given the tuned parameters for the simulator. An instance of the GALE system is configured and deployed to computational infrastructure.
  • The vehicle is deployed and piloted manually towards the operational area of the campaign.
  • The pilots execute the Mission Configuration Workflow (Section 5.D). This signals the start of the automated piloting, completing the setup for this deployment.
6 EXPERIMENTAL SETUP
6.A Field Experiments
@@
Our system was evaluated in two observational campaigns in the North Sea: MOGli (one glider) and eSWEETS3 (two gliders simultaneously). In these deployments, gliders were operated by NOC on behalf of external partners the UK Met Office and the University of East Anglia. A transect formed of two goal positions was used in all deployments. Deployment coordinates and summary statistics are given in Tables 1 and 2. A map of the three field deployments is shown in Figure 9.
Table 1: Deployment goal locations and parameters. All distances are given in km.
@@
Table 2: Deployment summary statistics.
@@
Figure 9: Field deployment locations in the North Sea.
MOGli. Met Office Glider (MOGli) is an ongoing joint project between the UK Met Office (MO) and NOC. Measurements of ocean properties are gathered by a single glider and transmitted periodically to the MO. This data is used to refine MO forecast models. The use of GALE to pilot the glider leads to a virtuous circle as AMM15 forecasts are used to derive navigation plans, while the scientific data itself is used to improve the forecasts.
The glider was deployed in a region with low bottom temperature variance since this was part of the MOGli scientific objectives. Fully autonomous control operated from 2 May to 3 July 2025, following preliminary experiments that began in November 2024. A pause in autonomous control was in place from 21 May to 7 June 2025 while the glider was recovered, recharged, and redeployed to the region.
The action space was chosen to contain relative bearings $\alpha\in\{-40,-20,0,20,40\}$. For the fixed control parameters $\mathbf{c}$, a value of $n_{\text{yos}}=5$ yos was chosen to lead to an operationally appropriate dive duration of approximately $3$ hours. A target depth of $z_{\text{bottom}}=150$ was used, which exceeds the depth of the seabed in this region. This was driven by the requirement to sample the entire water column, and hence the glider uses the altimeter to inflect when close to the seabed, a feature also supported by our simulator.
eSWEETS3. In the enabling Sustainable Wind Energy Expansion in Seasonally Stratified Seas (eSWEETS3) project, two Slocum gliders were deployed close to the Kincardine floating offshore wind farm. The gliders were positioned upstream (referred to as eSWEETS3 North) and downstream (eSWEETS3 South) of the wind farm to assess the impacts of the floating structures on the local turbulent mixing, which in turn influences oxygen availability, nutrient supplies, and biological productivity.
The project deployments were performed in two phases. In the first phase (May - August 2025), gliders were piloted manually using standard procedures. In the second phase (August - October 2025), GALE was used to control the gliders following preliminary experiments. Fully autonomous operation was in place from 2 September to 30 October 2025 for eSWEETS3 North and from 18 September to 30 October 2025 for eSWEETS3 South.
The chosen action space contained relative bearings $\alpha\in\{-90,-60,-30,0,30,60,90\}$, a wider range than in MOGli to account for the effect of strong tidal currents, which were aligned roughly orthogonal to the glider transects. $n_{\text{yos}}=10$ yos were used, leading to dives lasting approximately $4$ hours in this shallower region. A target depth of $z_{\text{bottom}}=150$ was used which, as in MOGli, exceeds the bathymetric limit.
@@
The eSWEETS3 deployments were subject to additional operational requirements. As the gliders were deployed in areas with strong tidal currents, there was a risk of vehicle loss if it drifted away from the mission area, especially into the wind farm or shipping lanes. Safety polygons were defined, and additional functionality was implemented such that, if the glider drifted outside of the safety polygon, the algorithm described in Section 5.B is replaced with a simplified navigation plan aiming directly for the center of the polygon. Normal navigation was resumed upon re-entry. Furthermore, as continuing to sample along the transect upon re-entry was deemed more important for the scientific objectives than turning back to achieve the goal, the system switched to the next goal if the glider had already progressed past 90% of the transect before exiting the safety polygon.
@@
Baseline. As a baseline for the Planner, we consider the Straight-to-Goal approach, which always selects the action with $\alpha=0$. For an equitable comparison, the glider remains under the autonomous control of GALE when Straight-to-Goal is used. Since only one of the two methods can control the glider at any one time, an alternating-control approach was taken. The two methods took turns to control the glider after the achievement of $2$ goals (equivalently, the algorithm was switched after every back-and-forth transect traversal). Even though the currents experienced by the two algorithms differ, we expect that any bias thus introduced is removed in expectation.
@@
For MOGli, alternating-control was followed throughout. For eSWEETS3, exclusive use of Straight-to-Goal was mandated by operational concerns until 18 September for eSWEETS3 North and 2 October for eSWEETS3 South. GALE was subsequently set to issue Planner waypoints for a comparable duration (until 14 October for eSWEETS3 North and 15 October for eSWEETS3 South). For the final period, the algorithms were switched to alternating-control after every back-and-forth transect traversal as in MOGli.
6.B Simulation Experiments
We present a set of experiments in simulation whose setup mirrors the deployments listed above. They are an idealised version of the real-world experiments, wherein file transfers always succeed and currents experienced are well-represented by the forecasts. Underwater goal achievement detection (Section 5.A) and the additional functionality for eSWEETS3 (Section 6.A) are not supported in simulation. The simulator and planner parameters are otherwise set identically to those used in the field deployments.
We select three scenarios for each deployment to be replayed in simulation out of the tens of transects completed in the real world (see counts in Table 2). A scenario consists of a starting state (glider position and time) and goal assigned to the glider, which implies a set of historical current forecasts corresponding to the starting state. The same scenarios can be ran as many times as desired with different realisations of the uncertainty, allowing statistically robust comparisons. The baseline remains the same (Straight-to-Goal). In contrast with the field experiments, in simulation the two methods can be compared using identical conditions.
To determine the scenarios, we first assess the number of dives required by Straight-to-Goal to complete each transect averaged over $200$ repetitions. We pick the starting conditions for the transects with the minimum (Favourable), median (Neutral), and maximum (Unfavourable) dives required on average for the baseline to reach the goal. This yields scenarios that represent a range of conditions experienced by the glider during the deployment periods. We then re-run both the Planner and Straight-to-Goal over $1000$ random seeds to report the final results.
6.C Metrics
@@
Planner. To evaluate the planner, we report the total transect duration (i.e., the cumulative cost as defined in Section 4.B), which is the objective that the method optimises for. As secondary metrics, we report the number of dives (highly correlated with total duration) and the total distance covered for each traversal. All ± entries in tables denote 95% confidence intervals. Occasionally, navigation plans were not generated or followed by the glider due to planned (e.g., software upgrades) or unplanned (e.g., anomalies causing abort of the dive) concerns. Even though these events typically impacted a single surfacing, they compromise the transect validity. Such transects were excluded from the planner evaluation.
@@
Simulator. To evaluate the simulator, we report the true and simulated dive distances and durations as well as their differences (deltas). Since datasets contain different numbers of dives, we report the mean (per-dive) log-likelihood $\bar{\ell}$ instead of the summed log-likelihood $\ell$. These metrics are assessed on the dives recorded during the live deployment, which were not used for simulator optimisation or validation. Furthermore, as the simulator is stochastic and produces several plausible next states, we report statistics with respect to the sample means obtained by drawing $40$ samples for each dive.
6.D Simulator and Planner Parameter Optimisation
@@
To optimise the simulator parameters (Section 4.D), we use historical glider data and current forecasts gathered in preliminary deployments (concretely, February - March 2025 for MOGli and May - June for eSWEETS3). The dataset for simulator parameter optimisation is pooled across the two eSWEETS3 gliders given the proximity of the two regions. Therefore, the same simulator parameters are used in both eSWEETS3 North and South. We use AMM15 forecasts archived by our system for simulator parameter optimisation. However, such data can also be obtained through publicly available oceanographic data repositories (e.g., [43] and [44]).
@@
The dives are split randomly into training and validation sets. Parameters are optimised using Bayesian optimisation (BO) on the training set, and the best parameters on the validation set with respect to $J(\Theta)$ are selected. We apply several data cleaning steps by excluding dives that did not terminate due to normal surfacing, those with speed $2$ standard deviations outside the mean, and with a difference between heading and surface location greater than a $75$ degree threshold for MOGli and $80$ for eSWEETS3. These criteria resulted in excluding $19.58\%$ of dives for MOGLi and $5.30\%$ dives for eSWEETS3. $n_{\text{SIM-BO}}=500$ iterations of BO were used as the scoring function plateaued before this step in all experiments.
@@
In terms of planner parameters, we use $n_{\text{threads}}=8$ parallel search trees for both deployments. We set $n_{\text{trials}}=10000$ and $n_{\text{trials}}=5000$ for MOGli and eSWEETS3 respectively, as for the latter deployment less data is transmitted and therefore less online planning time is available. We found that the planner hyperparameters (exploration constant, progressive widening parameters, maximum tree depth) are sensitive to simulator parameters. Therefore, after fixing simulator parameters, planner hyperparameters are optimised downstream with BO under the objective of minimising the cumulative cost to reach the goal.
7 RESULTS
@@
Table 3: Comparison of simulated and true dive metrics averaged over all dives in the real-world deployments.
@@
Simulated dive distance (m) 2119 4991 4362
Distance delta (m) 1448 1555 1080
@@
Duration delta (s) 345 1481 1377
Mean log-likelihood $\bar{\ell}$ -9.29 -78.67 -133.90
7.A Simulator Evaluation
Figure 10 shows the progress of the scoring function $J(\Theta)$ during the steps of the Bayesian optimisation procedure on the MOGli data. Additionally, in Figure 11, we show snapshots of simulator outputs for a MOGli dive at the beginning, intermediate, and final steps of the procedure. The refinement of the simulator output distribution with respect to the true surfacing location is evident.
Several values for the spread penalty $\lambda_{\text{reg}}$ were tested, and the parameters that obtained the best $J(\Theta)$ on the validation set were chosen ($\lambda_{\text{reg}}=1$ for MOGli, corresponding to the figures shown, and $\lambda_{\text{reg}}=0$ for eSWEETS3). The dive durations can realistically be treated as approximately constant in the MOGli region given the assumption of fixed control parameters and the fact that the bathymetry is consistent throughout the transect. Therefore, the duration score coefficient for MOGli is $\lambda_{\text{dur}}=0$. In contrast, we set $\lambda_{\text{dur}}=1$ in the eSWEETS3 region as the eastern ends of the eSWEETS3 transects are 204020-40 m deeper than the western ends. The variation in bathymetry induces variation in dive durations given fixed dive control parameters (i.e., dive durations increase in deeper regions). The standard deviation $\sigma=700$ was estimated based on data from the first phase which used manual piloting (Section 6.A).
Figure 11: Progress of the proposed simulator parameter optimisation procedure (Section 4.D) on one of the dives in the MOGli deployment. Simulator parameters are progressively adjusted such that the surfacing locations generated by the simulator (grey circles) best match the true surfacing location (red triangle).
In Table 3, we report several metrics concerning simulator performance on the deployment data (not seen during training or model selection). The distances of the dives output by the simulator were close to their true average on eSWEETS3, but noticeably underestimated for MOGli. The distance between the true and simulated surfacing locations is around $1.5$ km on average for all deployments (though it is inherently challenging to draw conclusions by comparing an individual sample with the empirical mean of the simulator distribution). In contrast, the dive durations were more accurate for MOGli given the aforementioned variation in bathymetry for eSWEETS3. The mean log-likelihood $\bar{\ell}$ degraded on the deployment data compared to the training data, as would be expected. On the deployment data, the optimised simulator parameters yielded scores of $\bar{\ell}=-9.29$ for MOGli and $\bar{\ell}=-106.00$ for eSWEETS3, which are significantly better than those of randomly chosen parameters at the start of optimisation ($\bar{\ell}\approx-32$ and $\bar{\ell}\approx-285$ respectively). This evidences the benefits of the proposed tuning procedure. The log-likelihood scores are worse for eSWEETS3 compared to MOGli due to the stronger currents in this tidal region.
We note that there are many possible causes for predictive errors in the simulator, which include the following: forecasts being wrong, the steady-state approximation of our simulator, mismatches between the physical glider used for simulator optimisation and the one used in the real world (distinct vehicles with variations in payload and ballasting), and noise in how the glider executes the control commands. It is highly likely that all these factors are at play and, while simulator accuracy can be improved, some degree of error is unavoidable.
7.B Planner Evaluation in Simulation
Table 4: Results comparing the Straight-to-Goal and Planner algorithms across several scenarios in simulation. Bold indicates the better-performing algorithm.
Deployment Scenario Algorithm Mean Duration (hrs) Mean Number Dives Mean Path Length (km)